1. The question is Write a linear equation for the cost of EACH band
2. Find the number of tickets for which the cost of the two bands will be equal.
Band A charges $600 to play for the entire evening.
Band B charges $350 plus $1.25 for each ticket sold.

Answer by stanbon(75887)

The question is Write a linear equation for the cost of EACH band
Band A charges $600 to play for the entire evening.
A(x) = $600
---
Band B charges $350 plus $1.25 for each ticket sold.
B(x) = 350 + 1.25x where x is the number of tickets sold.
----
2. Find the number of tickets for which the cost of the two bands will be equal.
A(x) = B(x)
600 = 350 + 1.25x
1.25x = 250
x = 200 (# of tickets sold to make the costs equal)
================
Cheers,
Stan H.